Assume compnay ABC faces the following total revenue and cost function TR=8Q and TC=Q2+4Q+2.find the maximum level of output and profit


1
Expert's answer
2022-02-11T08:32:44-0500

For profit maximization,

MR= MC

MR = 8

MC= 2Q+4

"\therefore2Q+4= 8"

"2Q= 4"

Q= 2

"\Pi= TR- TC"

8Q- ("Q^2+4Q+2)= 8\times2- (2^2+(4\times2)+2"

= 16-16= 0